# Delensing for Precision Cosmology: Optimizing Future CMB B-mode Surveys to Constrain $r$

### v2

@@ Abstract
The detection of primordial B-modes, a key probe of cosmic inflation, is increasingly challenged by contamination from weak gravitational lensing B-modes induced by large-scale structure (LSS). We present a delensing pipeline designed to enhance the sensitivity to the inflationary parameter $r$, minimizing reliance on foreground mitigation during lensing reconstruction. Using simulations of Simons Observatory-like CMB observations and Euclid-like LSS surveys in the Northern hemisphere, we demonstrate that excluding $-\ell$ modes ($\ell<200$) effectively mitigates foreground biases, enabling robust lensing potential reconstruction using observed CMB polarization maps. We reconstruct the lensing potential with a minimum-variance (MV) quadratic estimator (QE) applied to CMB polarization data and combine this with external LSS tracers to improve delensing efficiency. Two complementary methods—the Gradient-order template and the Inverse-lensing approach—are used to generate lensing B-mode templates, which are cross-correlated with observed B-modes. This achieves a 40% reduction in the uncertainty of r with CMB-only reconstruction, improving to 60% when incorporating external LSS tracers. We validate our results using both the Hamimeche & Lewis likelihood and a Gaussian approximation, finding consistent constraints on $r$. Our work establishes a streamlined framework for ground-based CMB experiments, demonstrating that synergies with LSS surveys significantly enhance sensitivity to primordial gravitational waves.
Keywords:
CMB polarization – CMB lensing – CMB delensing – Foreground – Large-scale structure
@@ 1 Introduction
One of the most significant challenges in achieving accurate measurements of CMB polarization is contamination from foreground emissions, particularly Galactic polarized foregrounds, which are highly complex and difficult to mitigate. Depending on the sky region, their amplitudes correspond roughly to $r$=0.01–0.1 at their minimum frequency range (70–90 GHz)krachmalnicoff2016characterization . Galactic polarized foregrounds are primarily composed of two components: synchrotron radiation, which dominates at lower frequencies (below 100 GHz), and thermal dust emission, which becomes significant at higher frequencies (above 100 GHz). Synchrotron radiation is produced by relativistic electrons spiraling in the Galactic magnetic field, while thermal dust emission arises from dust grains aligned with the same magnetic field. Fortunately, several component separation methods have been developed to disentangle foregrounds and the CMB based on their different frequency dependence, such as Commander eriksen2008joint ; adam2016planck , NILC basak2012needlet , SEVEM fernandez2012multiresolution , and SMICA cardoso2008component , among others. In this paper, we use the NILC method to examine the foreground components in the large aperture telescope (LAT) observations.
With the improvements in detector sensitivity, both current and future CMB experiments have enabled to measure the primordial fluctuations in the CMB B-mode polarization. The sensitivity is now such that contamination from lensing B-modes, rather than instrumental noise, is becoming the primary source of limitation. During the travel of CMB photons from the last scattering surface to us, they are gravitationally deflected by large-scale structures (LSS) in the universe. This gravitational lensing effect distorts the observed pattern of CMB anisotropies, with a typical deflection angle of about 2 arcminutes lewis2006weak . A portion of primordial E-modes is converted into B-modes during weak gravitational lensing, and the resulting B-modes contaminate measurements of primordial gravitational waves, making lensing B-mode a source of noise for primordial B-mode detection manzotti2017cmb . Therefore, reducing the contamination introduced by lensing becomes increasingly crucial when instrumental noise becomes subdominant compared to the lensed B-mode, which occurs at around 5 $\mu$K-arcmin.
@@ 2 CMB lensing and delensing
$\tilde{\Theta}(\hat{\mathbf{n}})$ $=\Theta(\hat{\mathbf{n}}+\mathbf{d}(\hat{\mathbf{n}}))$ (1)
@@
$[\tilde{Q}\pm i\tilde{U}](\hat{\mathbf{n}})$ $=[Q\pm iU](\hat{\mathbf{n}}+\mathbf{d}(\hat{\mathbf{n}}))$ (2)
@@
where tildes indicate lensed quantities and $\mathbf{d}$ is the deflection angle, given by the gradient of the lensing potential $\mathbf{d}(\hat{\mathbf{n}})=\nabla\phi(\hat{\mathbf{n}})$ (here we ignore the curl modes) namikawa2014lensing , which correlates to the lensing convergence as:
@@
×$\times\sqrt{\frac{(2\ell+1)(2\ell^{\prime}+1)(2L+1)}{16\pi}}\begin{pmatrix}\ell&\ell^{\prime}&0\\ -s&s^{\prime}&0\end{pmatrix}$
@@ 2.1 A brief introduction to the methods
The linear-order approximation (Eq. (3)) is known to be a good approximation for the lensing B-modes on large scales challinor2005lensed ; baleato2021limitations . Therefore, it is feasible to construct a lensing B-mode template on large scales using only gradient-order terms. This method is particularly convenient when working in the map domain, as shown by Eq. (2). In this case, the gradient-order template can be expressed as the product of the deflection angle and the gradients of the polarization fields , namely: $\nabla_{i}[Q\pm iU](\hat{\mathbf{n}})\nabla^{i}\phi(\hat{\mathbf{n}})$. In practice, the gradient of a field can be easily obtained using the ladder operators Okamoto:2003zw :
@@
where the primary CMB fields can be recovered from the lensed field by remapping the latter using the inverse deflection angle $\mathbf{d^{inv}}(\hat{\mathbf{n}})$:
$\tilde{X}(\hat{\mathbf{n}})=X(\hat{\mathbf{n}}+\mathbf{d}(\hat{\mathbf{n}}))\Leftrightarrow X(\hat{\mathbf{n}})=\tilde{X}(\hat{\mathbf{n}}+\mathbf{d^{inv}}(\hat{\mathbf{n}})).$ (7)
In practice, the inverse deflection angle is approximately the negative of the deflection at low-order approximation. A more accurate estimate can be obtained by solving Eq. (6) using a Newton-Raphson scheme, where the inverse deflection angle $\mathbf{d^{inv}}(\hat{\mathbf{n}})$ is iteratively computed as described in Carron:2017vfg . A lensing B-mode template can be obtained by subtracting the inverse-lensed B-mode map from the observed data.
@@ 2.2.1 Bias Analysis of the Gradient-Order Template Method
=$&\left\{\mathcal{B}^{(1)}[\mathcal{W}^{E}E^{\text{lens}}\ast\mathcal{W}^{\phi}\phi]\right\}+\mathcal{B}^{(1)}[\mathcal{W}^{E}E^{\text{noise}}\ast\mathcal{W}^{\phi}\phi]\\ &+\mathcal{B}^{(1)}[\mathcal{W}^{E}(E^{\text{lens}}+E^{\text{noise}})\ast\mathcal{W}^{\phi}\phi^{\text{noise}}]\\ &+\mathcal{B}^{(1)}[\mathcal{W}^{E}E^{\text{fg}}\ast\mathcal{W}^{\phi}(\phi+\phi^{\text{noise}})]\end{aligned}$
@@ 2.2.2 Bias analysis of inverse-lensing method
+$\quad+\left\{\mathcal{B}[\mathcal{W}^{E}B^{\text{lens}}\star\mathbf{d^{inv}_{\text{noise}}}]+\mathcal{B}[\mathcal{W}^{E}B^{\text{noise}}\star\mathbf{d^{inv}}]-\mathcal{W}^{E}B^{\text{lens}}\right\}$
@@ 3.2 Simulation of instrumental noise
27 GHz 91 arcmin 25 $\mu$K-arcmin (20,200) 7.4 arcmin 52 $\mu$K-arcmin \
39 GHz 63 arcmin 17 $\mu$K-arcmin (20,260) 5.1 arcmin 27 $\mu$K-arcmin \
93 GHz 30 arcmin 1.9 $\mu$K-arcmin (20,320) 2.2 arcmin 5.8 $\mu$K-arcmin (200,4800)
145 GHz 17 arcmin 2.1 $\mu$K-arcmin (20,600) 1.4 arcmin 6.3 $\mu$K-arcmin (200,6000)
225 GHz 11 arcmin 4.2 $\mu$K-arcmin (20,920) 1.0 arcmin 15 $\mu$K-arcmin \
280 GHz 9 arcmin 10 $\mu$K-arcmin (20,1100) 0.9 arcmin 37 $\mu$K-arcmin \
@@ 3.3 Simulation of foreground emission
To match actual foreground observation, we determine the parameters of the foreground model for the observed sky area (see Fig.5) using the foreground templates generated by PySM3 zonca2021python ; thorne2017python . We first use the d9s4 model krachmalnicoff2018s ; akrami2020planck in PySM3 to generate six foreground templates $F^{\text{Temp},EB}$ at 27 GHz, 39 GHz, 93 GHz, 145 GHz, 225 GHz, and 280 GHz, and calculate their power spectra $C_{\ell}^{\nu\nu^{\prime}}|_{FG}$. We then investigate the seven-parameter space $\{A_{d},\alpha_{d},\beta_{d},A_{s},\alpha_{s},\beta_{s},\epsilon_{DS}\}$ using Cobaya Torrado_2021 ; torrado2020cobaya . The peak values of the parameter posterior distribution are then considered a reasonable estimate of the true sky foreground. These values will be used in subsequent foreground simulations. Note that the spectral index $\beta_{c}$ should be $\beta_{c}(\hat{n})$ for strictly speaking, in this work we neglect spatial variations of the foreground SED and treat them as Gaussian random fields.
@@
$\alpha_{s}$0.800 $\mathcal{U}(-1.4,0)$
$\beta_{s}$3.100 $\mathcal{U}(-3.3,-2.7)$
@@ 3.4.1 LSS tracer models
which is just a constant at a fixed redshift. This is only a crude approximation and it is not self-consistent(e.g. $\delta_{g}(\hat{\mathbf{n}},z)$ will be smaller than $-1$ for $b>1$) and it is not preserved in time sigad2000measuring . It is reasonable to expect that, at any given time, scale, and for any galaxy type, galaxy biasing is generally nonlinear. In other words, the bias parameter $b_{\text{gal}}$ should vary as a function of $\delta_{m}$ and is typically stochastic, implying that a range of $\delta_{g}$ values may correspond to a given $\delta_{m}$. dekel1999stochastic proposed a comprehensive formalism for galaxy biasing that naturally distinguishes between nonlinearity and stochasticity. For more details, please refer to their work, as a detailed discussion is beyond the scope of this paper.
@@ 4.2.1 Internal lensing reconstruction
The reconstruction is performed strictly following the pipeline described in C. Note that for the polarization lensing reconstruction, we use the noise level accounted for the instrumental noise only in filtering, which introduces negligible sub-optimality since we have excluded the multipoles on large scales. We separately reconstruct the minimum-variance (MV) estimators $\hat{\phi}$ from the LAT observations at 93 GHz and 145 GHz for $6000200<\ell<$ 6000, with the maximum multipole for reconstruction set to $L_{\text{max}}=3071$. These are then combined using the inverse-variance weights:
@@
[…]
@@
Figure 8: Reconstructed lensing map. The left panel is the input lensing map; the middle panel is the Wiener-filtered reconstructed map from the MV quadratic estimator, the right panel is the difference between them. In order to highlight the lens structure, we plot the Wiener -filtered lensing deflection angle amplitude $\hat{\alpha}^{\text{WF}}=\sqrt{L(L+1)}\hat{\phi}_{LM}^{\text{MV}}C^{\phi\phi,\text{fid}}_{L}/[C^{\phi\phi,\text{fid}}_{L}+N^{(0),\text{ana}}_{L}]$, with multipole range $200020<L<$ 2000. Note that the edges of the map, which are not used in subsequent analysis, have been trimmed for clarity.
@@ 4.3 Lensed CMB maps processing
$\begin{gathered}Q_{\ell m}\Rightarrow\frac{C_{\ell}^{EE}}{C_{\ell}^{EE}+N_{\ell}^{EE}}Q_{\ell m},\\ U_{\ell m}\Rightarrow\frac{C_{\ell}^{EE}}{C_{\ell}^{EE}+N_{\ell}^{EE}}U_{\ell m}.\\ \end{gathered}$ (42)
@@ 4.4 Delensing
Regarding the Gradient-order lensing template method, we proceed by converting the spin-2 fields $Q+iU$ to spin-1 and spin-3 fields, and transform the spin-0 lensing potential field to spin-$-1$ and spin-1 fields using CMBlensplus namikawa2021CMBlensplus . It’s important to note that only E-modes should be included in the QU fields. Following this conversion, we compute the gradient lensing template according to Section 2.1. The resulting QU represents the lensing effect on E-mode, from which the lensing B-mode can be separated using harmonic transformation and EB leakage is corrected with an ”E-mode Recycling” method described in liu2019methods .
@@ 4.6.1 Baseline result
$C_{\ell}^{\text{SAT},\nu\nu^{\prime}}&=rC_{\ell}^{\text{tens}}|_{r=1}+A_{L}C_{\ell}^{\text{lens}}+N_{\ell}^{BB,\nu\nu^{\prime}}\\ &\quad+C_{\ell}^{\nu\nu^{\prime}}|_{\text{FG (7 params.)}},\end{aligned}$ (44)
@@
$N_{\ell}^{BB,\nu\nu^{\prime}}$ =$=\begin{cases}N_{\ell}^{\text{SAT},BB,\nu},&\text{if }\nu=\nu^{\prime}\\ 0,&\text{if }\nu\neq\nu^{\prime}.\end{cases}$ (47)
@@
$-2\ln\mathcal{L}(C_{\ell}|\hat{C}_{\ell})$ $=\frac{2l+1}{2}\mathrm{Tr}\left[\left(\mathbf{C}_{f_{\ell}}^{-1/2}\mathbf{C}_{g_{\ell}}\mathbf{C}_{f_{\ell}}^{-1/2}\right)^{2}\right]$ (49)
@@
Gradient-order Method Inverse-lensing Method Parameter Input value Before adding LT After adding LT (Internal) After adding LT (Combined tracer) After adding LT (Internal) After adding LT (Combined tracer) $r(\times 10^{3})$ 0 $-0.055\pm 1.800$ $-0.038\pm 0.958$ $-0.030\pm 0.649$ $-0.033\pm 0.984$ $-0.028\pm 0.660$ $A_{L}$ 1 $1.000\pm 0.049$ $1.000\pm 0.042$ $1.000\pm 0.037$ $1.000\pm 0.042$ $1.000\pm 0.037$ $A_{d}(\mu K^{2})$ 14.300 $14.215\pm 0.640$ $14.215\pm 0.637$ $14.222\pm 0.633$ $14.219\pm 0.638$ $14.223\pm 0.633$ $\alpha_{d}$0.650 $-0.642\pm 0.073$ $-0.642\pm 0.072$ $-0.643\pm 0.072$ $-0.642\pm 0.072$ $-0.643\pm 0.072$ $\beta_{d}$ 1.48 $1.480\pm 0.010$ $1.480\pm 0.009$ $1.480\pm 0.007$ $1.480\pm 0.009$ $1.480\pm 0.008$ $A_{s}(\mu K^{2})$ 2.400 $2.378\pm 0.175$ $2.378\pm 0.170$ $2.378\pm 0.167$ $2.378\pm 0.170$ $2.379\pm 0.168$ $\alpha_{s}$0.800 $-0.775\pm 0.157$ $-0.778\pm 0.152$ $-0.778\pm 0.149$ $-0.778\pm 0.153$ $-0.778\pm 0.149$ $\beta_{s}$3.100 $-3.102\pm 0.035$ $-3.101\pm 0.031$ $-3.101\pm 0.028$ $-3.101\pm 0.031$ $-3.101\pm 0.028$ $\epsilon_{ds}(\times 10^{2})$ 0 $0.003\pm 3.811$ $0.003\pm 3.729$ $0.012\pm 3.684$ $0.019\pm 3.730$ $0.016\pm 3.678$
@@
Gradient-order Method Inverse-lensing Method Parameter Input value Before adding LT After adding LT (Internal) After adding LT (Combined tracer) After adding LT (Internal) After adding LT (Combined tracer) $r(\times 10^{3})$ 0 $0.259\pm 1.829$ $0.287\pm 0.984$ $0.178\pm 0.668$ $0.295\pm 1.010$ $0.182\pm 0.680$ $A_{L}$ 1 $1.001\pm 0.049$ $1.000\pm 0.042$ $1.001\pm 0.037$ $1.000\pm 0.042$ $1.001\pm 0.037$ $A_{d}(\mu K^{2})$ 14.300 $14.392\pm 0.645$ $14.379\pm 0.640$ $14.382\pm 0.638$ $14.380\pm 0.640$ $14.381\pm 0.637$ $\alpha_{d}$0.650 $-0.656\pm 0.072$ $-0.655\pm 0.072$ $-0.655\pm 0.072$ $-0.655\pm 0.072$ $-0.655\pm 0.071$ $\beta_{d}$ 1.48 $1.480\pm 0.010$ $1.480\pm 0.008$ $1.480\pm 0.007$ $1.480\pm 0.008$ $1.480\pm 0.007$ $A_{s}(\mu K^{2})$ 2.400 $2.421\pm 0.175$ $2.419\pm 0.171$ $2.420\pm 0.168$ $2.420\pm 0.172$ $2.419\pm 0.168$ $\alpha_{s}$0.800 $-0.816\pm 0.154$ $-0.814\pm 0.149$ $-0.814\pm 0.146$ $-0.815\pm 0.150$ $-0.815\pm 0.147$ $\beta_{s}$3.100 $-3.102\pm 0.034$ $-3.102\pm 0.031$ $-3.102\pm 0.028$ $-3.102\pm 0.031$ $-3.102\pm 0.028$ $\epsilon_{ds}(\times 10^{2})$ 0 $-0.020\pm 3.781$ $-0.031\pm 3.705$ $-0.009\pm 3.656$ $-0.020\pm 3.699$ $-0.017\pm 3.639$

### v3

@@ Abstract
The detection of primordial B-modes, a key probe of cosmic inflation, is increasingly challenged by contamination from weak gravitational lensing B-modes induced by large-scale structure (LSS). We present a delensing pipeline designed to enhance the sensitivity to the inflationary parameter $r$, minimizing reliance on foreground mitigation during lensing reconstruction. Using simulations of Simons Observatory-like CMB observations and Euclid-like LSS surveys in the Northern hemisphere, we demonstrate that excluding low-$\ell$ modes ($\ell<200$) effectively mitigates foreground biases, enabling robust lensing potential reconstruction using observed CMB polarization maps. We reconstruct the lensing potential with a minimum-variance (MV) quadratic estimator (QE) applied to CMB polarization data and combine this with external LSS tracers to improve delensing efficiency. Two complementary methods—the Gradient-order template and the Inverse-lensing approach—are used to generate lensing B-mode templates, which are cross-correlated with observed B-modes. This achieves a 40% reduction in the uncertainty of r with CMB-only reconstruction, improving to 60% when incorporating external LSS tracers. We validate our results using both the Hamimeche & Lewis likelihood and a Gaussian approximation, finding consistent constraints on $r$. Our work establishes a streamlined framework for ground-based CMB experiments, demonstrating that synergies with LSS surveys significantly enhance sensitivity to primordial gravitational waves.
@@ 1 Introduction
One of the most significant challenges in achieving accurate measurements of CMB polarization is contamination from foreground emissions, particularly Galactic polarized foregrounds, which are highly complex and difficult to mitigate. Depending on the sky region, their amplitudes correspond roughly to rr=0.01–0.1 at their minimum frequency range (70–90 GHz)krachmalnicoff2016characterization . Galactic polarized foregrounds are primarily composed of two components: synchrotron radiation, which dominates at lower frequencies (below 100 GHz), and thermal dust emission, which becomes significant at higher frequencies (above 100 GHz). Synchrotron radiation is produced by relativistic electrons spiraling in the Galactic magnetic field, while thermal dust emission arises from dust grains aligned with the same magnetic field. Fortunately, several component separation methods have been developed to disentangle foregrounds and the CMB based on their different frequency dependence, such as Commander eriksen2008joint ; adam2016planck , NILC basak2012needlet , SEVEM fernandez2012multiresolution , and SMICA cardoso2008component , among others. In this paper, we use the NILC method to examine the foreground components in the large aperture telescope (LAT) observations.
[…]
@@ 2 CMB lensing and delensing
[…]
@@
[…]
@@
[…]
@@
$\times\sqrt{\frac{(2\ell+1)(2\ell^{\prime}+1)(2L+1)}{16\pi}}\begin{pmatrix}\ell&\ell^{\prime}&0\\ -s&s^{\prime}&0\end{pmatrix}$
@@ 2.1 A brief introduction to the methods
[…]
@@
[…]
$).\tilde{X}(\hat{\mathbf{n}})=X(\hat{\mathbf{n}}+\mathbf{d}(\hat{\mathbf{n}}))\Leftrightarrow X(\hat{\mathbf{n}})=\tilde{X}(\hat{\mathbf{n}}+\mathbf{d^{inv}}(\hat{\mathbf{n}}$)). (7)
[…]
@@ 2.2.1 Bias Analysis of the Gradient-Order Template Method
$=\begin{aligned} &\left\{\mathcal{B}^{(1)}[\mathcal{W}^{E}E^{\text{lens}}\ast\mathcal{W}^{\phi}\phi]\right\}+\mathcal{B}^{(1)}[\mathcal{W}^{E}E^{\text{noise}}\ast\mathcal{W}^{\phi}\phi]\\ &+\mathcal{B}^{(1)}[\mathcal{W}^{E}(E^{\text{lens}}+E^{\text{noise}})\ast\mathcal{W}^{\phi}\phi^{\text{noise}}]\\ &+\mathcal{B}^{(1)}[\mathcal{W}^{E}E^{\text{fg}}\ast\mathcal{W}^{\phi}(\phi+\phi^{\text{noise}})]\end{aligned}$
@@ 2.2.2 Bias analysis of inverse-lensing method
$\quad+\left\{\mathcal{B}[\mathcal{W}^{E}B^{\text{lens}}\star\mathbf{d^{inv}_{\text{noise}}}]+\mathcal{B}[\mathcal{W}^{E}B^{\text{noise}}\star\mathbf{d^{inv}}]-\mathcal{W}^{E}B^{\text{lens}}\right\}$
@@ 3.2 Simulation of instrumental noise
[…]
@@ 3.3 Simulation of foreground emission
[…]
@@
$\alpha_{s}$ -0.800 $\mathcal{U}(-1.4,0)$
$\beta_{s}$ -3.100 $\mathcal{U}(-3.3,-2.7)$
@@ 3.4.1 LSS tracer models
which is just a constant at a fixed redshift. This is only a crude approximation and it is not self-consistent(e.g. $\delta_{g}(\hat{\mathbf{n}},z)$ will be smaller than 1-1 for $b>1$) and it is not preserved in time sigad2000measuring . It is reasonable to expect that, at any given time, scale, and for any galaxy type, galaxy biasing is generally nonlinear. In other words, the bias parameter $b_{\text{gal}}$ should vary as a function of $\delta_{m}$ and is typically stochastic, implying that a range of $\delta_{g}$ values may correspond to a given $\delta_{m}$. dekel1999stochastic proposed a comprehensive formalism for galaxy biasing that naturally distinguishes between nonlinearity and stochasticity. For more details, please refer to their work, as a detailed discussion is beyond the scope of this paper.
@@ 4.2.1 Internal lensing reconstruction
The reconstruction is performed strictly following the pipeline described in C. Note that for the polarization lensing reconstruction, we use the noise level accounted for the instrumental noise only in filtering, which introduces negligible sub-optimality since we have excluded the multipoles on large scales. We separately reconstruct the minimum-variance (MV) estimators $\hat{\phi}$ from the LAT observations at 93 GHz and 145 GHz for $200<\ell<6000$, with the maximum multipole for reconstruction set to $L_{\text{max}}=3071$. These are then combined using the inverse-variance weights:
@@
Here, $\hat{\phi}^{\text{MV},\nu_{i}}_{LM}$ represents the MV estimator derived from the observed maps at frequency $\nu_{i}$ (where $\nu_{i}$ $\in$ $\{93\text{GHz}$, $145\text{GHz}\}$), and $N^{(0),\nu_{i}}_{L}$ denotes the corresponding reconstruction noise power spectrum.
@@
Figure 8: Reconstructed lensing map. The left panel is the input lensing map; the middle panel is the Wiener-filtered reconstructed map from the MV quadratic estimator, the right panel is the difference between them. In order to highlight the lens structure, we plot the Wiener -filtered lensing deflection angle amplitude $\hat{\alpha}^{\text{WF}}=\sqrt{L(L+1)}\hat{\phi}_{LM}^{\text{MV}}C^{\phi\phi,\text{fid}}_{L}/[C^{\phi\phi,\text{fid}}_{L}+N^{(0),\text{ana}}_{L}]$, with multipole range $20<L<2000$. Note that the edges of the map, which are not used in subsequent analysis, have been trimmed for clarity.
@@ 4.3 Lensed CMB maps processing
[…]
@@ 4.4 Delensing
Regarding the Gradient-order lensing template method, we proceed by converting the spin-2 fields $Q+iU$ to spin-1 and spin-3 fields, and transform the spin-0 lensing potential field to spin-1-1 and spin-1 fields using CMBlensplus namikawa2021CMBlensplus . It’s important to note that only E-modes should be included in the QU fields. Following this conversion, we compute the gradient lensing template according to Section 2.1. The resulting QU represents the lensing effect on E-mode, from which the lensing B-mode can be separated using harmonic transformation and EB leakage is corrected with an ”E-mode Recycling” method described in liu2019methods .
@@ 4.6.1 Baseline result
$\begin{aligned} C_{\ell}^{\text{SAT},\nu\nu^{\prime}}&=rC_{\ell}^{\text{tens}}|_{r=1}+A_{L}C_{\ell}^{\text{lens}}+N_{\ell}^{BB,\nu\nu^{\prime}}\\ &\quad+C_{\ell}^{\nu\nu^{\prime}}|_{\text{FG (7 params.)}},\end{aligned}$ (44)
@@
$N_{\ell}^{BB,\nu\nu^{\prime}}$ $=\begin{cases}N_{\ell}^{\text{SAT},BB,\nu},&\text{if }\nu=\nu^{\prime}\\ 0,&\text{if }\nu\neq\nu^{\prime}.\end{cases}$ (47)
@@
[…]
@@
Gradient-order Method Inverse-lensing Method Parameter Input value Before adding LT After adding LT (Internal) After adding LT (Combined tracer) After adding LT (Internal) After adding LT (Combined tracer) $r(\times 10^{3})$ 0 $055\pm 1.800 0.0$$38\pm 0.958 0.0$$0\pm 0.649 0.033$$3\pm 0.984 0.0$$028\pm 0.660$ $A_{L}$ 1 $1.000\pm 0.049$ $1.000\pm 0.042$ $1.000\pm 0.037$ $1.000\pm 0.042$ $1.000\pm 0.037$ $A_{d}(\mu K^{2})$ 14.300 $14.215\pm 0.640$ $14.215\pm 0.637$ $14.222\pm 0.633$ $14.219\pm 0.638$ $14.223\pm 0.633$ $\alpha_{d}$ -0.650 $0.642\pm 0.073$ $0.642\pm 0.072$ $0.643\pm 0.072$ $0.642\pm 0.072$ $0.643\pm 0.072$ $\beta_{d}$ 1.48 $1.480\pm 0.010$ $1.480\pm 0.009$ $1.480\pm 0.007$ $1.480\pm 0.009$ $1.480\pm 0.008$ $A_{s}(\mu K^{2})$ 2.400 $2.378\pm 0.175$ $2.378\pm 0.170$ $2.378\pm 0.167$ $2.378\pm 0.170$ $2.379\pm 0.168$ $\alpha_{s}$ -0.800 $0.775\pm 0.157$ $0.778\pm 0.152$ $0.778\pm 0.149$ $0.778\pm 0.153$ $0.778\pm 0.149$ $\beta_{s}$ -3.100 $3.102\pm 0.035$ $3.101\pm 0.031$ $3.101\pm 0.028$ $3.101\pm 0.031$ $3.101\pm 0.028$ $\epsilon_{ds}(\times 10^{2})$ 0 $003\pm 3.811 0.0$$03\pm 3.729 0.0$$12\pm 3.684 0.0$$19\pm 3.730 0.0$$016\pm 3.678$
@@
Gradient-order Method Inverse-lensing Method Parameter Input value Before adding LT After adding LT (Internal) After adding LT (Combined tracer) After adding LT (Internal) After adding LT (Combined tracer) $r(\times 10^{3})$ 0 $0.259\pm 1.829$ $0.287\pm 0.984$ $0.178\pm 0.668$ $0.295\pm 1.010$ $0.182\pm 0.680$ $A_{L}$ 1 $1.001\pm 0.049$ $0\pm 0.042 1.001±0.0$$1\pm 0.037$ $0\pm 0.042 1.001±0.0$$1\pm 0.037$ $A_{d}(\mu K^{2})$ 14.300 $14.392\pm 0.645$ $14.379\pm 0.640$ $14.382\pm 0.638$ $14.380\pm 0.640$ $14.381\pm 0.637$ $\alpha_{d}$ -0.650 $0.656\pm 0.072$ $0.655\pm 0.072$ $0.655\pm 0.072$ $0.655\pm 0.072$ $0.655\pm 0.071$ $\beta_{d}$ 1.48 $1.480\pm 0.010$ $1.480\pm 0.008$ $1.480\pm 0.007$ $1.480\pm 0.008$ $1.480\pm 0.007$ $A_{s}(\mu K^{2})$ 2.400 $2.421\pm 0.175$ $2.419\pm 0.171$ $2.420\pm 0.168$ $2.420\pm 0.172$ $2.419\pm 0.168$ $\alpha_{s}$ -0.800 $0.816\pm 0.154$ $0.814\pm 0.149$ $0.814\pm 0.146$ $0.815\pm 0.150$ $0.815\pm 0.147$ $\beta_{s}$ -3.100 $3.102\pm 0.034$ $3.102\pm 0.031$ $3.102\pm 0.028$ $3.102\pm 0.031$ $3.102\pm 0.028$ $\epsilon_{ds}(\times 10^{2})$ 0 $020\pm 3.781 0.0$$31\pm 3.705 0.0$$09\pm 3.656 0.0$$20\pm 3.699 0.0$$017\pm 3.639$